OpenStudy (anonymous):

Which shows solutions to x + 3y = 7 if x and y must be whole numbers {(0, 5), (1, 3), (2, 1)} {(2, 1), (1, 4), (0, 7)} {(1, 2), (4, 1), (7, 0)} {(0, 5), (1, 4), (2, 1)}

OpenStudy (anonymous):

Plug each pair in for x and y and see which ones work...ie which ones give you 7. In each pair (__,__) plug the first number in for x and the 2nd for y

jimthompson5910 (jim_thompson5910):

plug each point into the equation and see if you get a true result for instance, plug in (0,5) to get x + 3y = 7 0 + 3*5 = 7 .. plug in x = 0, y = 5 0 + 15 = 7 15 = 7 ... this is FALSE so because the last equation is false, this means that (0,5) is NOT a solution to x+3y = 7

OpenStudy (anonymous):

ok so this answer would be the {(0, 5), (1, 4), (2, 1)}??

jimthompson5910 (jim_thompson5910):

how can the answer be that if I just proved that (0,5) is not a solution to x+3y = 7

OpenStudy (anonymous):

oh sph sorry my mind isblank right now

OpenStudy (anonymous):

what if I think the answer is #2 is that right

jimthompson5910 (jim_thompson5910):

pick any point from that set and plug it into the equation

jimthompson5910 (jim_thompson5910):

if you get a true result, then check the other points if you get a false result, then you can eliminate that option

OpenStudy (anonymous):

well can you just tell me if I got that answer right? #2?

jimthompson5910 (jim_thompson5910):

let's plug in (2,1) x + 3y = 7 2 + 3*1 = 7 2 + 3 = 7 5 = 7 that's false, so (2,1) is NOT a solution

jimthompson5910 (jim_thompson5910):

see how this works?

OpenStudy (anonymous):

lol kinda of ok well what about #1

jimthompson5910 (jim_thompson5910):

(0,5) is part of choice A, so what do you think?

jimthompson5910 (jim_thompson5910):

remember above (0,5) is shown to not be a solution to x+3y = 7

OpenStudy (anonymous):

nope so I imam say its 3 right

jimthompson5910 (jim_thompson5910):

choice C is the only thing left, but I recommend you check all the points to make sure

OpenStudy (anonymous):

what this is a mid untit test I needa know is it right...

jimthompson5910 (jim_thompson5910):

you can't ask for help on tests